Solve the following linear equations.  x/2 – 1/5 = x/3 + ¼ 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

x/2–1/5=x/3+1⁄4
⇒x/2–x/3=1⁄4+1/5
⇒(3x–2x)/6=(5+4)/20
⇒3x–2x=9/20×6
⇒x=54/20
⇒x=27/10

Thanks